OpenStudy (anonymous):

what is the remainder when f(x)=-9x^17+x^13-x^2+8x is divided by x+1?

OpenStudy (anonymous):

f(-1) = -9(-1)^17+(-1)^13-(-1)^2+8(-1) = 9-1-1-8=-1 by remainder theorem, the remainder is -1

OpenStudy (ash2326):

We have \[ f(x)=-9x^{17}+x^{13}-x^2+8x\] We have to divide it by x+1 Use remainder theorem here put x=-1 in f(x) which will give us the remainder \[f(-1)=-9(-1)^{17}+(-1)^{-13} -(-1)^2+8(-1)=-9 \times -1+ (-1) -(1)-8\] we get \[f(-1)=9-1-1-8=-1\] so the remainder is -1

OpenStudy (anonymous):

oh ok thank you